OpenStudy (anonymous):

Algebra 2, Rational Functions... Simplify. Identify any x-values for which the expression is undefined. (problems below) medal for best help

OpenStudy (anonymous):

\[\frac{ 4x ^{6} }{ 2x ^{4} }\]

OpenStudy (anonymous):

rule of exponent: x^m/x^n=x^(m-n) so with that mind,